# A Prompt Intention Framework for Complex Workflow Generation

### v1

@@ 1 Introduction
The Opus Prompt Intention Framework is based on the following concepts introduced in Opus: A Large Workflow Model for Complex Workflow Generation by Fagnoni et al. [1] and Opus: A Workflow Intention Framework for Complex Workflow Generation by Kingston et al. [2]:
@@ 2 Background
We adopt the Opus Workflow Framework (Fagnoni et al. [1]) where a Workflow is represented as a Directed Acyclic Graph (DAG) consisting of Input Nodes, Task Nodes, Output Nodes, and Edges that define the Task execution flow from Inputs to Outputs. We generate Workflows at a semantic level: each Node is described using structured semantic features following a predefined schema, typically represented as a JSON object with fields such as name, description, and steps for Task Nodes. The schema is flexible and can be adapted as required, demonstrating the extensibility of the framework. Additionally, we adopt the Opus Workflow Intention Framework (Kingston et al. [2]), which maps business artefacts and Intention Elicitation into Workflow Signals—specifically Input (), Process (), and Output () elements- and into an Intention Set. In this work, we focus exclusively on Intention Elicitation as user queries, in text format. Since we operate directly at a semantic level using LLMs, each Workflow Signal is represented as a list of strings. This aligns with the Opus Workflow Intention Framework, where Workflow Signals are modeled as vectors and classified against generative families of Input, Process, and Output string elements. In our representation, each string corresponds to a distinct semantic component, enforced through prompt design. The generation of these strings is unconstrained beyond this granularity. Although further classification against generative families is possible, we do not consider this step essential to the core argument of this paper. Intention objects, forming an Intention Set for each Intention Elicitation, are represented as triples , where each element is a list of strings.
@@
The core argument of this paper is that Workflow Generation is significantly improved when guided by explicit Workflow Intentions. To substantiate this claim, we evaluate generated Workflows using two categories of metrics: (1) semantic and structural metrics, which assess Workflow similarity, and (2) LLM-based metrics, where LLMs are prompted to evaluate Workflow quality under predefined criteria. For semantic and structural metrics, we adopt metrics used in the Opus Workflow Framework [1]. These include BLEU (Bilingual Evaluation Understudy, Papineni et al. [22]), a precision-based metric widely used in machine translation; ROUGE (Recall-Oriented Understudy for Gisting Evaluation, Lin et al. [23]) and its variants—ROUGE-N (N-gram overlap), ROUGE-L (Longest common subsequence), and ROUGE-S (Skip-bigram co-occurrence)—commonly applied in summarization evaluation; METEOR (Metric for Evaluation of Translation with Explicit ORdering, Banerjee et al. [24]), which extends BLEU by incorporating stemming, synonym matching, and a higher recall weighting; and BERTScore (Zhang et al. [25]), which leverages contextual embeddings from pre-trained language models (specifically BERT) to measure textual similarity. For LLM-based evaluation, we adopt the LLM-as-a-Judge framework, where LLMs are used to evaluate model outputs. Jiawei et al. [26] provide a comprehensive overview of this approach, addressing challenges related to bias, robustness, and alignment with human judgments. When properly configured, LLM-based evaluation provides a powerful tool for qualitative assessment of complex Workflows. However, the inherent imprecision of LLM judgments—stemming from variability in model outputs and susceptibility to bias—requires mitigation strategies such as prompt calibration and repeated sampling. Employing multiple models as evaluators can help temper these limitations by diversifying judgment perspectives.
@@ 3 System Overview
We implement a reproducible Intention Capture system using LLMs, designed to be model-agnostic and demonstrated that even a basic Intention Capture improves Workflow Generation unequivocally. The system consists of two LLMs, as illustrated in Figure 1: the first extracts Input, Process, and Output Signals from an Intention Elicitation ; the second generates the Intention Set, where each Intention object comprises aligned Input, Process, and Output Signals. In this context, a Signal is a list of granular string elements within its category (Input, Process, or Output), mirroring the classification of encoded artifacts into predefined elements of families as defined in the Opus Workflow Intention Framework [2]. We evaluate the accuracy of the generated Signals and Intention Set using loss functions analogous to those employed in the Opus Workflow Intention Framework.
@@ 4 Data and Methodology
For each Intention , we generate a corresponding Intention Elicitation (as a user query) - a string crafted from , ensuring these components are explicitly stated and incorporated. To increase linguistic entropy and make the mapping non-injective, we introduce controlled variations in phrasing that preserve semantic content. This simulates user query diversity—akin to adding noise to a signal without loosing information. From , we generate a Workflow W, represented as a string (structure detailed in Appendix A.1, Figure 6). To simulate multi-intent scenarios, we sample a variable number of Intention Objects from to form an Intention Set, comprising the Reference Intention Sets:
@@
Intention Generation: Compute the loss between Decoded Signals and the Reference aggregated Signals:
@@ 4.1 Signal Extraction
where denotes the set of all functions mapping to .
@@ 4.2 Intention Generation
For Intention Generation , we prompt an LLM to generate an Intention Set, composed of Workflow Intention objects. The prompt is conditioned on both the Intention Elicitation and the extracted Workflow Signals. Each Intention object consists of three lists: (Input), (Process) and (Output). Let denotes the Decoded (LLM-generated) Intention Set, being the Reference Intention Set. We extend the Signal Loss to define the Intention Loss, which measures the alignment between the Decoded and Reference Intention Sets. The Intention Loss is defined as:
where denotes the set of all injections from to . The injection constraint ensures that each Decoded Intention can only be matched to a unique Reference Intention, preventing multiple predictions from mapping to the same ground-truth Intention.
@@ 4.3 Intention for Workflow Generation
The distance is a monotonically increasing function: as the similarity between two Workflows increases, grows accordingly. Given two sets of Workflows , we define the aggregated set-level Workflow Similarity as:
This formulation finds the optimal pairing between Decoded and Reference Workflows to maximize total Similarity. The normalization by the cardinality of the smaller set penalizes missing or redundant Workflows. We employ multiple Workflow Similarity metrics, reporting the square root of certain squared metrics (such as squared Euclidean distances or squared losses), converting them back to their original units and improving comparability with standard similarity measures.
@@ 6 Results
This section presents the results of Signal Extraction, Intention Generation, and Workflow Generation, both with and without Intention guidance. We conduct experiments using 9 leading LLMs across 10 Mixed Intention Levels. Each Level corresponds to the number of Complete Intentions combined within an Intention Elicitation, simulating query complexity. For each Level , 100 samples are generated, each representing a unique Intention Elicitation constructed from Complete Intentions. Unless otherwise specified, all experiments include the following models: OpenAI’s GPT-4.5, GPT-4o, o3-mini, o1, Anthropic’s Claude 3.5 Opus, Claude 3.5 Sonnet, Claude 3.7 Sonnet, DeepSeek V3, and R1.
@@ A.4 Workflow Generation Evaluation by LLM
A.4.2 LLM-as-a-Judge Scores - Claude 3.7 Sonnet
@@ A.4.3 Semantic and Structural Metrics - OpenAI o1
A.4.4 LLM-as-a-Judge Scores - OpenAI o1
@@ A.4.5 Semantic and Structural Metrics - OpenAI o3-mini
A.4.6 LLM-as-a-Judge Scores - OpenAI o3-mini
@@
A.4.8 LLM-as-a-Judge Scores - OpenAI GPT-4.5
@@ A.4.9 Semantic and Structural Metrics - OpenAI GPT-4o
A.4.10 LLM-as-a-Judge Scores - OpenAI GPT-4o
@@
A.4.12 LLM-as-a-Judge Scores - Claude 3.5 Sonnet
@@ A.4.13 Semantic and Structural Metrics - DeepSeek-V3
A.4.14 LLM-as-a-Judge Scores - DeepSeek-V3
@@ A.4.15 Semantic and Structural Metrics - DeepSeek-R1
A.4.16 LLM-as-a-Judge Scores - DeepSeek-R1

### v2

@@ 1 Introduction
The Opus Prompt Intention Framework is based on the following concepts introduced in Opus: A Large Work Model for Complex Workflow Generation by Fagnoni et al. [1] and Opus: A Workflow Intention Framework for Complex Workflow Generation by Kingston et al. [2]:
@@ 2 Background
We adopt the Opus Workflow Framework (Fagnoni et al. [1]) where a Workflow is represented as a Directed Acyclic Graph (DAG) consisting of Input Nodes, Task Nodes, Output Nodes, and Edges that define the Task execution flow from Inputs to Outputs. We generate Workflows at a semantic level: each Node is described using structured semantic features following a predefined schema, typically represented as a JSON object with fields such as name, description, and steps for Task Nodes. The schema is flexible and can be adapted as required, demonstrating the extensibility of the framework. Additionally, we adopt the Opus Workflow Intention Framework (Kingston et al. [2]), which maps business artefacts and Intention Elicitation into Workflow Signals—specifically Input (), Process (), and Output () elements—and into an Intention Set. In this work, we focus exclusively on Intention Elicitation as user queries, in text format. Since we operate directly at a semantic level using LLMs, each Workflow Signal is represented as a list of strings. This aligns with the Opus Workflow Intention Framework, where Workflow Signals are modeled as vectors and classified against generative families of Input, Process, and Output string elements. In our representation, each string corresponds to a distinct semantic component, enforced through prompt design. The generation of these strings is unconstrained beyond this granularity. Although further classification against generative families is possible, we do not consider this step essential to the core argument of this paper. Intention objects, forming an Intention Set for each Intention Elicitation, are represented as triples , where each element is a list of strings.
@@
The core argument of this paper is that Workflow Generation is significantly improved when guided by explicit Workflow Intentions. To substantiate this claim, we evaluate generated Workflows using two categories of metrics: (1) semantic and structural metrics, which assess Workflow similarity, and (2) LLM-based metrics, where LLMs are prompted to evaluate Workflow quality under predefined criteria. For semantic and structural metrics, we adopt metrics used in the Opus Workflow Framework [1]. These include BLEU (Bilingual Evaluation Understudy, Papineni et al. [22]), a precision-based metric widely used in machine translation; ROUGE (Recall-Oriented Understudy for Gisting Evaluation, Lin et al. [23]) and its variants—ROUGE-N (N-gram overlap), ROUGE-L (Longest common subsequence), and ROUGE-S (Skip-bigram co-occurrence)—commonly applied in summarization evaluation; METEOR (Metric for Evaluation of Translation with Explicit ORdering, Banerjee et al. [24]), which extends BLEU by incorporating stemming, synonym matching, and a higher recall weighting; and BERTScore (Zhang et al. [25]), which leverages contextual embeddings from pre-trained language models (specifically BERT) to measure textual similarity. For LLM-based evaluation, we adopt the LLM-as-a-Judge framework, where LLMs are used to evaluate model outputs. Gu et al. [26] provide a comprehensive overview of this approach, addressing challenges related to bias, robustness, and alignment with human judgments. When properly configured, LLM-based evaluation provides a powerful tool for qualitative assessment of complex Workflows. However, the inherent imprecision of LLM judgments—stemming from variability in model outputs and susceptibility to bias—requires mitigation strategies such as prompt calibration and repeated sampling. Employing multiple models as evaluators can help temper these limitations by diversifying judgment perspectives.
@@ 3 System Overview
We implement a reproducible Intention Capture system using LLMs, designed to be model-agnostic and demonstrate that even a basic Intention Capture improves Workflow Generation unequivocally. The system consists of two LLMs, as illustrated in Figure 1: the first extracts Input, Process, and Output Signals from an Intention Elicitation; the second generates the Intention Set, where each Intention object comprises aligned Input, Process, and Output Signals. In this context, a Signal is a list of granular string elements within its category (Input, Process, or Output), mirroring the classification of encoded artefacts into predefined elements of families as defined in the Opus Workflow Intention Framework [2]. We evaluate the accuracy of the generated Signals and Intention Set using loss functions analogous to those employed in the Opus Workflow Intention Framework.
@@ 4 Data and Methodology
For each Intention , we generate a corresponding Intention Elicitation (as a user query) —a string crafted from , ensuring these components are explicitly stated and incorporated. To increase linguistic entropy and make the mapping non-injective, we introduce controlled variations in phrasing that preserve semantic content. This simulates user query diversity—akin to adding noise to a signal without losing information. From , we generate a Workflow W, represented as a string (structure detailed in Appendix A.1, Figure 6). To simulate multi-intent scenarios, we sample a variable number of Intention Objects from to form an Intention Set, comprising the Reference Intention Sets:
@@
Signal Loss: Compute the loss between Decoded Signals and the Reference aggregated Signals:
@@ 4.1 Signal Extraction
where denote the set of all functions mapping to .
@@ 4.2 Intention Generation
For Intention Generation , we prompt an LLM to generate an Intention Set, composed of Workflow Intention objects. The prompt is conditioned on both the Intention Elicitation and the extracted Workflow Signals. Each Intention object consists of three lists: (Input), (Process) and (Output). Let denote the Decoded (LLM-generated) Intention Set, being the Reference Intention Set. We extend the Signal Loss to define the Intention Loss, which measures the alignment between the Decoded and Reference Intention Sets. The Intention Loss is defined as:
where denote the set of all injections from to . The injection constraint ensures that each Decoded Intention can only be matched to a unique Reference Intention, preventing multiple predictions from mapping to the same ground-truth Intention.
@@ 4.3 Intention for Workflow Generation
The similarity is a monotonically increasing function: as the similarity between two Workflows increases, grows accordingly. Given two sets of Workflows , we define the aggregated set-level Workflow Similarity as:
This formulation finds the optimal pairing between Decoded and Reference Workflows to maximize total Similarity. The normalization by the cardinality of the larger set penalizes missing or redundant Workflows. We employ multiple Workflow Similarity metrics, reporting the square root and scaled version of certain metrics (such as Euclidean distances or squared losses), converting them back to their original units and improving comparability with standard similarity measures.
@@ 6 Results
This section presents the results of Signal Extraction, Intention Generation, and Workflow Generation, both with and without Intention guidance. We conduct experiments using 9 leading LLMs across 10 Mixed Intention Levels. Each Level corresponds to the number of Complete Intentions combined within an Intention Elicitation, simulating query complexity. For each Level , 100 samples are generated, each representing a unique Intention Elicitation constructed from Complete Intentions. Unless otherwise specified, all experiments include the following models: OpenAI’s GPT-4.5, GPT-4o, o3-mini, o1, Anthropic’s Claude 3 Opus, Claude 3.5 Sonnet, Claude 3.7 Sonnet, DeepSeek V3, and R1.
@@ A.4 Workflow Generation Evaluation by LLM
A.4.2 LLM-as-a-Judge Scores (scaled to [0,1]) - Claude 3.7 Sonnet
@@ A.4.3 Semantic and Structural Metrics - OpenAI o1
A.4.4 LLM-as-a-Judge Scores (scaled to [0,1]) - OpenAI o1
@@ A.4.5 Semantic and Structural Metrics - OpenAI o3-mini
A.4.6 LLM-as-a-Judge Scores (scaled to [0,1]) - OpenAI o3-mini
@@
A.4.8 LLM-as-a-Judge Scores (scaled to [0,1]) - OpenAI GPT-4.5
@@ A.4.9 Semantic and Structural Metrics - OpenAI GPT-4o
A.4.10 LLM-as-a-Judge Scores (scaled to [0,1]) - OpenAI GPT-4o
@@
A.4.12 LLM-as-a-Judge Scores (scaled to [0,1]) - Claude 3.5 Sonnet
@@ A.4.13 Semantic and Structural Metrics - DeepSeek-V3
A.4.14 LLM-as-a-Judge Scores (scaled to [0,1]) - DeepSeek-V3
@@ A.4.15 Semantic and Structural Metrics - DeepSeek-R1
A.4.16 LLM-as-a-Judge Scores (scaled to [0,1]) - DeepSeek-R1